OpenStudy (anonymous):

Amy and Stephanie were playing a number game. Amy picked a number and told Stephanie that if she added 8 to the number and multiplied the sum by 2, it would equal 6. Which equation could Stephanie use to find the number, x? @apoorvk

OpenStudy (apoorvk):

Okay let me assume the number picked by Amy be 'x'. Cool? So, when Steph adds 8 to it, and then multiplies the whole thing by 2, what would it look like?

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

Thanks @e.mccormick

OpenStudy (anonymous):

@Abhisar

OpenStudy (abhisar):

I am gonna guide you make this equation up..right?

OpenStudy (anonymous):

Yesh

OpenStudy (anonymous):

I hate direct answers.

OpenStudy (abhisar):

So, let's add 8 to that number x. How would you write that?

OpenStudy (abhisar):

#Waiting for your Response!

OpenStudy (anonymous):

#SORRYcomeback

OpenStudy (anonymous):

@Abhisar

OpenStudy (abhisar):

#StillWaitingForResponse

OpenStudy (anonymous):

Please no hashtags, gosh geez

OpenStudy (anonymous):

And idk how to write equations

OpenStudy (abhisar):

Haven't you learned it?

OpenStudy (abhisar):

Then how will you solve this question :-(

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

8(x + 2) = 6 i think this is it.

OpenStudy (abhisar):

Nerps, You need to first add 8 to x, Step1: 8+x

OpenStudy (anonymous):

okay

OpenStudy (abhisar):

Now you need to multiply the whole sum with 2. STEP 2: Use brackets and multiply (8+x) with 2 ----> How will you write this?

OpenStudy (anonymous):

2 + (x + 8) = 6 8(x + 2) = 6 2(x + 8) = 6 2(x) + 8 = 6 these are the answers